the length of the hypotenuse of an isosceles right triangle is 8.2 centimeters. find the perimeter of the triangle to nearest tenth.
Answer by stanbon(57361)

the length of the hypotenuse of an isosceles right triangle is 8.2 centimeters. find the perimeter of the triangle to nearest tenth.
---------------------------------
isosceles implies the legs are equal.
Let each leg be "x".
x^2 + x^2 = 8.2^2
2x^2 = 67.24
x^2 = 33.62
x = 5.798 cm
------------------
Perimeter = 2(5.798) + 8.2
P = 19.8 cm
============================
Cheers,
Stan H.
============================